# Exact inference via quasi-conjugacy in two-parameter Poisson–Dirichlet hidden Markov models

### v1

@@ 13 $r\leftarrow r-X_{i}$
1 Draw $W\sim\mathrm{Beta}(\theta+\alpha l(\pi),|\pi|-\alpha l(\pi))$
2 Draw $X^{(1)}\sim\mathrm{Dir}(\pi_{1}-\alpha,\dots,\pi_{l(\pi)}-\alpha)$
3 Draw $X^{(2)}\sim\mathrm{PD}_{\alpha,\theta+\alpha l(\pi)}$ as in Algorithm 1 with truncation error $\varepsilon/(1-W)$
@@ 7     Initialization: $s_{\mu}\leftarrow 0$ if $s_{\mu}$ is NULL
Input: Partitions $\pi^{0:N}$; weights $\{w_{\lambda},\lambda\in\Lambda\}$ from $\mathrm{Pr}_{\alpha,\theta}(X(t)\in{\rm d}x|\pi^{0:N})$; partition size $m$
@@ 1 Draw $\lambda\sim w_{\lambda}$
A.5.2 Weights of filtering and smoothing distributions
@@
$\Pr(\Pi_{1:|\lambda|}=\lambda,\Pi_{|\lambda|+1:|\lambda|+|\omega|}=\omega)=\sum_{\mu\in\mathrm{coag}(\lambda,\omega)}\mathcal{H}(\lambda,\omega\mid\mu)\,\mathrm{PSF}_{\alpha,\theta}(\mu),$ (89)
@@
$\Pr(\Pi_{1:|\lambda|+|\omega|}=\mu\mid\Pi_{1:|\lambda|}=\lambda,\Pi_{|\lambda|+1:|\lambda|+|\omega|}=\omega)=\frac{\mathcal{H}(\lambda,\omega\mid\mu)\,\mathrm{PSF}_{\alpha,\theta}(\mu)}{\sum_{\mu^{\prime}\in\mathrm{coag}(\lambda,\omega)}\mathcal{H}(\lambda,\omega\mid\mu^{\prime})\,\mathrm{PSF}_{\alpha,\theta}(\mu^{\prime})}.$ (90)
@@
$s_{\mu}(\Delta_{k+1})\ \propto\ \sum_{\lambda\in\Lambda_{0:k}}\sum_{\omega\in\Omega^{k:N}}\frac{v_{\lambda}}{\mathrm{PSF}_{\alpha,\theta}(\lambda)}\,\frac{h_{\omega}(\Delta_{k+1})}{\mathrm{PSF}_{\alpha,\theta}(\omega)}\,\mathcal{H}(\lambda,\omega\mid\mu)\,\mathrm{PSF}_{\alpha,\theta}(\mu),$ (91)
@@
A.5.3 Further figures
Figure 12 displays filtering (left column) and smoothing (right column) estimates of the three largest system coordinates (top to bottom), as described in Sections A.4.1 and A.4.2. The plotted data (diamonds) are the observed relative frequencies at the corresponding ranks. For example, the second panel shows the relative frequency of the second-largest group. Although presented separately for visualization, these estimates are obtained jointly for the full $\nabla$-valued vector, and the data are likewise used jointly to condition the model.

### v2

@@ 13 $r\leftarrow r-X_{i}$
1 Draw $W\sim\mathrm{Beta}(\theta+\ell\alpha,|\pi|-\ell\alpha)$
2 Draw $V^{(1)}\sim\mathrm{Dir}(\pi_{1}-\alpha,\dots,\pi_{\ell}-\alpha)$
3 Draw $V^{(2)}\sim\mathrm{PD}_{\alpha,\theta+\alpha\ell}$ as in Algorithm 1 with truncation error $\varepsilon/(1-W)$
@@ 7     Initialization: $s_{\mu}\leftarrow 0$ if $s_{\mu}$ is NULL
Input: Partitions $\pi^{0:N}$; weights $\{w_{\lambda},\lambda\in\Lambda\}$ from $\Pr(X(t)\in{\rm d}x|\pi^{0:N})$; partition size $m$
@@ 1 Draw $\lambda\sim w_{\lambda}$
A.4.2 Weights of filtering and smoothing distributions
@@
$\Pr(\Pi_{1:|\lambda|}=\lambda,\Pi_{|\lambda|+1:|\lambda|+|\omega|}=\omega)=\sum_{\mu\in\mathrm{coag}(\lambda,\omega)}\mathcal{H}(\lambda,\omega|\mu)\,\mathrm{PSF}_{\alpha,\theta}(\mu),$ (92)
@@
$\Pr(\Pi_{1:|\lambda|+|\omega|}=\mu|\Pi_{1:|\lambda|}=\lambda,\Pi_{|\lambda|+1:|\lambda|+|\omega|}=\omega)=\frac{\mathcal{H}(\lambda,\omega|\mu)\,\mathrm{PSF}_{\alpha,\theta}(\mu)}{\sum_{\mu^{\prime}\in\mathrm{coag}(\lambda,\omega)}\mathcal{H}(\lambda,\omega|\mu^{\prime})\,\mathrm{PSF}_{\alpha,\theta}(\mu^{\prime})}.$ (93)
@@
$s_{\mu}(\Delta_{k+1})\ \propto\ \sum_{\lambda\in\Lambda_{0:k}}\sum_{\omega\in\Omega^{k:N}}\frac{v_{\lambda}}{\mathrm{PSF}_{\alpha,\theta}(\lambda)}\,\frac{h_{\omega}(\Delta_{k+1})}{\mathrm{PSF}_{\alpha,\theta}(\omega)}\,\mathcal{H}(\lambda,\omega|\mu)\,\mathrm{PSF}_{\alpha,\theta}(\mu),$ (94)
@@
A.4.3 Further figures
Figure 12 displays filtering (left column) and smoothing (right column) estimates of the three largest system coordinates (top to bottom), as described in Sections A.3.1 and A.3.2. The plotted data (diamonds) are the observed relative frequencies at the corresponding ranks. For example, the second panel shows the relative frequency of the second-largest group. Although presented separately for visualization, these estimates are obtained jointly for the full $\nabla$-valued vector, and the data are likewise used jointly to condition the model.